# Interpreting CLIP with Hierarchical Sparse Autoencoders

### v1

@@ Abstract
Sparse autoencoders (SAEs) are useful for detecting and steering interpretable features in neural networks, with particular potential for understanding complex multimodal representations. Given their ability to uncover interpretable features, SAEs are particularly valuable for analyzing large-scale vision-language models (e.g., CLIP and SigLIP), which are fundamental building blocks in modern systems yet remain challenging to interpret and control. However, current SAE methods are limited by optimizing both reconstruction quality and sparsity simultaneously, as they rely on either activation suppression or rigid sparsity constraints. To this end, we introduce Matryoshka SAE (MSAE), a new architecture that learns hierarchical representations at multiple granularities simultaneously, enabling a direct optimization of both metrics without compromise. MSAE establishes a new state-of-the-art Pareto frontier between reconstruction quality and sparsity for CLIP, achieving 0.99 cosine similarity and less than 0.1 fraction of variance unexplained while maintaining  80% sparsity. Finally, we demonstrate the utility of MSAE as a tool for interpreting and controlling CLIP by extracting over 120 semantic concepts from its representation to perform concept-based similarity search and bias analysis in downstream tasks like CelebA.
@@
Figure 1: Refer to caption Matryoshka Sparse Autoencoder (MSAE) enables learning hierarchical concept representations from coarse to fine-grained features while avoiding rigid sparsity constraints in TopK and the activation shrinkage problem in ReLU SAE. (B) At training, MSAE uses multiple top-$k$ values up to dimension $d$ instead of a single $k$ like in TopK SAE, combining losses across different granularities. (C) At inference our method uses the whole $d$-dimensional representation. (D) MSAE allows for more precise editing and manipulation in the concept space.
@@ 1 Introduction
Despite CLIP’s widespread adoption, understanding how it processes and represents information remains a challenge. The distributed nature of its learned representations and the complexity of the optimized loss function make it particularly difficult to interpret. Traditional explainability approaches have limited success in addressing this challenge: gradient-based feature attributions (Simonyan, 2013; Shrikumar et al., 2017; Selvaraju et al., 2017; Sundararajan et al., 2017; Abnar & Zuidema, 2020) struggle to provide human-interpretable explanations, perturbation-based approaches (Zeiler & Fergus, 2014; Ribeiro et al., 2016; Lundberg, 2017; Adebayo et al., 2018) yield inconsistent results, and concept-based methods (Ramaswamy et al., 2023; Oikarinen et al., 2023) are constrained by their reliance on manually curated concept datasets. This interpretability gap hinders our ability to identify and mitigate potential biases or failure modes of CLIP in downstream applications. Recent advances in mechanical interpretability (Conmy et al., 2023; Bereska & Gavves, 2024) use sparse autoencoders (SAEs) as a tool for disentangling interpretable features in neural networks (Cunningham et al., 2024). When applied to CLIP’s representation space, SAEs offer the potential to decompose complex, distributed representations into human-interpretable components through self-supervised learning. It eliminates the need for concept datasets and limits predefined concept sets in favor of natural concept emergence.
@@
However, training effective SAEs poses unique challenges. The richness of data distribution and high dimensionality of CLIP’s multimodal embedding space requires tuning the sparsity-reconstruction trade-off (Bricken et al., 2023; Gao et al., 2024). Furthermore, evaluating SAE effectiveness extends beyond traditional metrics, requiring discovering interpretable features that maintain their semantic meaning across both visual and textual modalities. Current approaches for enforcing sparsity in autoencoders use either $L_{1}$ (Bricken et al., 2023) or TopK (Gao et al., 2024) proxy functions, each with significant drawbacks. $L_{1}$ regularization results in activation shrinkage, systematically underestimating feature activations and potentially missing subtle but important concepts. TopK enforces a fixed number of active neurons, imposing rigid constraints that may not align with the natural concept density in different regions of CLIP’s embedding space (Gao et al., 2024; Bussmann et al., 2024b).
@@
To this end, we propose a hierarchical approach to sparse autoencoders, a new architecture inspired by Matryoshka representation learning ([Uncaptioned image], Kusupati et al., 2022), as illustrated in Figure 1. While Matryoshka SAE (MSAE) can be applied to interpret any neural network representation, we demonstrate its utility in the CLIP’s complex multimodal embedding space. At its core, MSAE applies TopK operations $h$-times with progressively increasing numbers of $k$ neurons, learning representations at $h$ granularities simultaneously – from coarse concepts to fine-grained features. By combining reconstruction losses across all granularity levels, MSAE achieves a more flexible and adaptive sparsity pattern. We effectively remove the rigid constraints of simple TopK while avoiding the activation shrinkage problems associated with $L_{1}$ regularization, resulting in the state-of-the-art Pareto frontier between the reconstruction quality and sparsity.
@@ 2 Related Work
Mechanistic interpretability. Mechanistic interpretability seeks to reverse engineer neural networks analogously to decompiling computer programs (Conmy et al., 2023; Bereska & Gavves, 2024). While early approaches focus on generating natural language descriptions of individual neurons (Hernandez et al., 2021; Bills et al., 2023), the polysemantic nature of neural representations makes this challenging. A breakthrough comes with sparse autoencoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2024), which demonstrate the ability to recover monosemantic features. Recent architectural advancements like Gated (Rajamanoharan et al., 2024) and TopK SAE variants (Gao et al., 2024) improve the sparsity–reconstruction trade-off, enabling successful application to LLMs (Templeton et al., 2024), diffusion models (Surkov et al., 2024), and medical imaging (Abdulaal et al., 2024). Recent work on SAE-based interpretation of CLIP embeddings (Rao et al., 2024) shows promise in extracting interpretable features.
Concept-based explainability. Concept-based explanations provide interpretability by identifying human-coherent concepts within neural networks’ latent spaces. While early approaches relied on manually curated concept datasets (Kim et al., 2018; Zhou et al., 2018), recent work has explored automated concept extraction (Ghorbani et al., 2019) and explicit concept learning (Liu et al., 2020; Koh et al., 2020; Espinosa Zarlenga et al., 2022), with successful applications in out-of-distribution detection (Madeira et al., 2023), image generation (Misino et al., 2022), and medicine (Lucieri et al., 2020). However, existing methods often struggle to scale to modern transformer architectures with hundreds of millions of parameters. Our approach addresses this limitation by first training SAE without supervision on concept learning, then efficiently mapping unit-norm decoder columns to defined vocabulary concepts using cosine similarity with CLIP embeddings.
@@ 3.1 Preliminaries
Existing approaches established two primary sparsity mechanisms. ReLU SAE (Bricken et al., 2023) uses $L_{1}$ regularization with the objective $\mathcal{L}(x):=\|x-\hat{x}\|_{2}^{2}+\lambda\|z\|_{1}$, while TopK SAE (Gao et al., 2024) enforces fixed sparsity through $z=\mathrm{ReLU}\left(\mathrm{TopK}\left(W_{\mathrm{enc}}\left(x-b_{\mathrm{pre}}\right)+b_{\mathrm{enc}}\right)\right)$. However, each approach faces distinct limitations: $L_{1}$ regularization causes activation shrinkage (Rajamanoharan et al., 2024), while TopK imposes rigid sparsity constraints (Bussmann et al., 2024b).
@@ 3.2 Matryoshka SAE Architecture
Sparsity coefficient weighting. We propose and evaluate two strategies for setting the weighting coefficients $\alpha_{i}$. The uniform weighting (UW) approach sets $\alpha_{i}=1$ for all $i$, while the reverse weighting (RW) strategy uses $\alpha_{i}=h-i+1$, giving higher weights to lower $k$ values. RW places greater emphasis on the sparsity of the representation as compared to UW. RW still maintains the quality of the reconstruction, as evidenced by our empirical findings of improved sparsity without significant performance degradation (Table 1).
@@ 4 Evaluating MSAE
In this section, we conduct extensive experiments to evaluate MSAE against ReLU and TopK SAEs. We compare the sparsity–fidelity trade-off (Section 4.2), at multiple granularity levels  (Section 4.3). We follow with evaluating the semantic quality of learned representations beyond traditional distance metrics (Section 4.4), analyzing decoder orthogonality (Section 4.5), and examining the statistical properties of SAE activation magnitudes (Section 4.6). To verify that MSAE successfully learns hierarchical features, we conduct experiments on the progressive recovery task (Section 4.7). We conclude with an ablation study comparing the influence of different training modalities in Section 4.8.
@@
Table 1: Quantitative comparison of SAE models on ImageNet-1k. We compare the following SAEs with expansion rate 8: ReLU with varying sparsity regularization ($\lambda$), TopK with 64 or 256 active neurons, and Matryoshka using uniform (UW) or reverse weighting (RW) $\alpha$ coefficients. Arrows ($\uparrow$/$\downarrow$) indicate the preferred direction of metrics. NDN values in parentheses show the dead neuron count on the training set. LP (KL) values are scaled by $10^{6}$ for readability. Extended results for higher expansion rates and the text modality are reported in Appendix F.5.
@@ 4.1 Evaluation Metrics
$L_{0}$ denotes the mean proportion of zero elements in SAE activations. Fraction of variance unexplained (FVU), also known as Normalized MSE (Gao et al., 2024), measures reconstruction fidelity by normalizing the mean squared reconstruction error $\mathcal{L}(x)$ by the mean squared value of the (mean-centered) input. Explained variance ratio (EVR) is FVU’s complement metric, defined as $1-\text{FVU}$. Linear probing (LP) assesses how well SAE preserves semantic information in the reconstructed embeddings on the downstream task. To evaluate this, we train a linear probe model on ImageNet-1k using CLIP embeddings as a backbone, with AdamW optimizer ($\mathrm{lr}=1e{-}3$), ReduceLROnPlateau scheduler, and batch_size=256. We measure performance by comparing predictions from original versus reconstructed embeddings using two metrics: Kullback-Leibler divergence (KL) between predicted class distributions and classification accuracy (Acc), where accuracy uses $\mathrm{argmax}$ predictions from original embeddings as targets. Centered kernel nearest neighbor alignment (CKNNA) (Huh et al., 2024) measures kernel alignment based on mutual nearest neighbors, providing a quantitative assessment of alignment between SAE activations and input embeddings. A detailed explanation is provided in Appendix E. Decoder orthogonality (DO) calculates the mean cosine similarity of the lower triangular portion of the SAE decoder, where 0 indicates perfect orthogonality. This metric assesses how orthogonal the monosemantic feature directions are in the decoder. Number of dead neurons (NDN) is a metric that measures how many neurons remain consistently inactive (zero in the SAE activations layer) across all inputs during training or evaluation, indicating the network’s inability to fully utilize its capacity for learning semantic features.
Figure 2: Comparison of sparsity–fidelity trade-offs across SAE architectures on ImageNet-1k. Each model presents results from all 3 expansion rates, comparing ReLU SAE ($\lambda=\{0.001,0.003,0.03\}$), TopK SAE ($k=\{32,64,128,256\}$), and MSAE (UW, RW). The optimal SAE would occupy the upper right corner, achieving both high sparsity and reconstruction fidelity. For extended results across both modalities, refer to Figure 11.
@@ 4.2 Sparsity–Fidelity Trade-off
We assess SAE performance using sparsity–fidelity trade-off, measuring sparsity with $L_{0}$ and reconstruction quality with EVR, following previous work. Figure 2 reveals that ReLU SAE shows difficulty balancing performance, achieving either high fidelity with low $L_{0}$ or the opposite, with expansion rate primarily improving sparsity. TopK SAE with higher k values achieves better but not ReLU-level fidelity while offering improved sparsity yet consistently suffering from at least 5% of dead neurons (Table 1). Both variants of MSAE achieve better sparsity than ReLU and better fidelity than TopK, establishing a superior Pareto frontier while maintaining less than 1% of dead neurons. The RW variant further improves sparsity as expected, with only minor fidelity degradation. Notably, only Matryoshka consistently improves on both metrics with higher expansion rates, while TopK struggles with increased expansion rate, and ReLU shows improvements only in the highest $\lambda$.

### v2

@@ Abstract
Sparse autoencoders (SAEs) are useful for detecting and steering interpretable features in neural networks, with particular potential for understanding complex multimodal representations. Given their ability to uncover interpretable features, SAEs are particularly valuable for analyzing vision-language models (e.g., CLIP and SigLIP), which are fundamental building blocks in modern large-scale systems yet remain challenging to interpret and control. However, current SAE methods are limited by optimizing both reconstruction quality and sparsity simultaneously, as they rely on either activation suppression or rigid sparsity constraints. To this end, we introduce Matryoshka SAE (MSAE), a new architecture that learns hierarchical representations at multiple granularities simultaneously, enabling a direct optimization of both metrics without compromise. MSAE establishes a state-of-the-art Pareto frontier between reconstruction quality and sparsity for CLIP, achieving 0.99 cosine similarity and less than 0.1 fraction of variance unexplained while maintaining  80% sparsity. Finally, we demonstrate the utility of MSAE as a tool for interpreting and controlling CLIP by extracting over 120 semantic concepts from its representation to perform concept-based similarity search and bias analysis in downstream tasks like CelebA. We make the codebase available at https://github.com/WolodjaZ/MSAE.
@@
Figure 1: Refer to caption Matryoshka Sparse Autoencoder (MSAE) enables learning hierarchical concept representations from coarse to fine-grained features while avoiding rigid sparsity constraints in TopK and the activation shrinkage problem in ReLU SAE. (B) At training, MSAE uses multiple top-$k$ values up to dimension $d$ instead of a single $k$ like in TopK SAE, combining losses across different granularities. (C) At inference, our method uses the whole $d$-dimensional representation. (D) MSAE allows for more precise editing and manipulation in the concept space.
@@ 1 Introduction
Despite CLIP’s widespread adoption, understanding how it processes and represents information remains a challenge. The distributed nature of its learned representations and the complexity of the optimized loss function make it particularly difficult to interpret. Traditional explainability approaches have limited success in addressing this challenge: gradient-based feature attributions (Simonyan, 2013; Shrikumar et al., 2017; Selvaraju et al., 2017; Sundararajan et al., 2017; Abnar & Zuidema, 2020) struggle to provide human-interpretable explanations, perturbation-based approaches (Zeiler & Fergus, 2014; Ribeiro et al., 2016; Lundberg & Lee, 2017; Adebayo et al., 2018; Baniecki et al., 2025) yield inconsistent results, and concept-based methods (Ramaswamy et al., 2023; Oikarinen et al., 2023) are constrained by their reliance on manually curated concept datasets. This interpretability gap hinders our ability to identify and mitigate potential biases or failure modes of CLIP in downstream applications (Biecek & Samek, 2024). Recent advances in mechanical interpretability (Conmy et al., 2023; Bereska & Gavves, 2024) use sparse autoencoders (SAEs) as a tool for disentangling interpretable features in neural networks (Cunningham et al., 2024). When applied to CLIP’s representation space, SAEs offer the potential to decompose complex, distributed representations into human-interpretable components through self-supervised learning. It eliminates the need for concept datasets and limits predefined concept sets in favor of natural concept emergence.
@@
However, training effective SAEs poses unique challenges. The richness of data distribution and high dimensionality of CLIP’s multimodal embedding space require tuning the sparsity-reconstruction trade-off (Bricken et al., 2023; Gao et al., 2025). Furthermore, evaluating SAE effectiveness extends beyond traditional metrics, requiring the discovery of interpretable features that maintain their semantic meaning across both visual and textual modalities. Current approaches for enforcing sparsity in autoencoders use either $L_{1}$ (Bricken et al., 2023) or TopK (Gao et al., 2025) proxy functions, each with significant drawbacks. $L_{1}$ regularization results in activation shrinkage, systematically underestimating feature activations and potentially missing subtle but important concepts. TopK enforces a fixed number of active neurons, imposing rigid constraints that may not align with the natural concept density in different regions of CLIP’s embedding space (Gao et al., 2025; Bussmann et al., 2024).
@@
To this end, we propose a hierarchical approach to sparse autoencoders, a new architecture inspired by Matryoshka representation learning ([Uncaptioned image], Kusupati et al., 2022), as illustrated in Figure 1. While Matryoshka SAE (MSAE) can be applied to interpret any neural network representation, we demonstrate its utility in the CLIP’s complex multimodal embedding space. At its core, MSAE applies TopK operations $h$-times with progressively increasing numbers of $k$ neurons, learning representations at $h$ granularities simultaneously – from coarse concepts to fine-grained features. By combining reconstruction losses across all granularity levels, MSAE achieves a more flexible and adaptive sparsity pattern. We remove the rigid constraints of simple TopK while avoiding the activation shrinkage problems associated with $L_{1}$ regularization, resulting in the state-of-the-art Pareto frontier between the reconstruction quality and sparsity.
@@ 2 Related Work
Mechanistic interpretability. Mechanistic interpretability seeks to reverse engineer neural networks analogously to decompiling computer programs (Conmy et al., 2023; Bereska & Gavves, 2024). While early approaches focus on generating natural language descriptions of individual neurons (Hernandez et al., 2021; Bills et al., 2023), the polysemantic nature of neural representations makes this challenging. A breakthrough comes with sparse autoencoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2024), which demonstrate the ability to recover monosemantic features. Recent architectural advancements like Gated (Rajamanoharan et al., 2024a) and TopK SAE variants (Gao et al., 2025) improve the sparsity–reconstruction trade-off, enabling successful application to LLMs (Templeton et al., 2024), diffusion models (Surkov et al., 2024), and medical imaging (Abdulaal et al., 2024). Recent work on SAE-based interpretation of CLIP embeddings (Rao et al., 2024) shows promise in extracting interpretable features.
Concept-based explainability. Concept-based explanations provide interpretability by identifying human-coherent concepts within neural networks’ latent spaces. While early approaches relied on manually curated concept datasets (Kim et al., 2018; Zhou et al., 2018; Bykov et al., 2023), recent work has explored automated concept extraction (Ghorbani et al., 2019; Kopf et al., 2024) and explicit concept learning (Liu et al., 2020; Koh et al., 2020; Espinosa Zarlenga et al., 2022), with successful applications in out-of-distribution detection (Madeira et al., 2023), image generation (Misino et al., 2022), and medicine (Lucieri et al., 2020). However, existing methods often struggle to scale to modern transformer architectures with hundreds of millions of parameters. Our approach addresses this limitation by first training SAE without supervision on concept learning, then efficiently mapping unit-norm decoder columns to defined vocabulary concepts using cosine similarity with CLIP embeddings.
@@ 3.1 Preliminaries
Existing approaches established two primary sparsity mechanisms. ReLU SAE (Bricken et al., 2023) uses $L_{1}$ regularization with the objective $\mathcal{L}(x):=\|x-\hat{x}\|_{2}^{2}+\lambda\|z\|_{1}$, while TopK SAE (Gao et al., 2025) enforces fixed sparsity through $z=\mathrm{ReLU}\left(\mathrm{TopK}\left(W_{\mathrm{enc}}\left(x-b_{\mathrm{pre}}\right)+b_{\mathrm{enc}}\right)\right)$. However, each approach faces distinct limitations: $L_{1}$ regularization causes activation shrinkage (Rajamanoharan et al., 2024a), while TopK imposes rigid sparsity constraints. A recent effort to address the rigidity of TopK is BatchTopK (Bussmann et al., 2024), which replaces the standard $\mathrm{TopK}$ function with $\mathrm{BatchTopK}$ within the TopK SAE method. The $\mathrm{BatchTopK}$ function treats all batch activations as a single, flattened vector before applying $\mathrm{TopK}$. This allows for a flexible number of active features per sample, with the total number of active features across the batch averaging to $k\times\text{batch size}$. Although BatchTopK relaxes the fixed sparsity of traditional TopK, it still relies on a predetermined $k$ parameter that requires careful tuning and continues to suffer from the potential for certain features to become ‘dead’ or rarely activated if they consistently fall outside the TopK selection.
@@ 3.2 Matryoshka SAE Architecture
Sparsity coefficient weighting. We propose and evaluate two strategies for setting the weighting coefficients $\alpha_{i}$. The uniform weighting (UW) approach sets $\alpha_{i}=1$ for all $i$, while the reverse weighting (RW) strategy uses $\alpha_{i}=h-i+1$, giving higher weights to lower $k$ values. By weighting the loss more heavily for sparser reconstructions, RW encourages the model to learn features that maintain reconstruction quality at lower k values. As shown in Table 1, this results in improved sparsity without significant performance degradation as the model learns that sparse representations achieve better loss even with slightly worse reconstruction quality, compared to UW which focuses primarily on reconstruction quality.
3.3 Training and Inference
CLIP embeddings exhibit misalignment across modalities, which can impact SAE training convergence and cross-modal transferability. Following (Bhalla et al., 2024), we normalize embeddings to ensure consistent behavior across modalities. We first center embeddings by subtracting the per-modality mean estimated from the training dataset. Next, we scale the centered embeddings by a dataset-computed scaling factor following (Conerly et al., 2024) to obtain $\mathbb{E}_{x\in\mathcal{X}}[\|x\|_{2}]=\sqrt{n}$. This scaling ensures that $\lambda$ has consistent effects across different CLIP architectures and modalities. For training, we compute the mean vector and scaling factor from the image modality. During inference on text embeddings, we apply the text-specific mean and scaling factor. Additionally, at inference, we remove TopK constraints from TopK-trained models, allowing the model to adaptively select the number of active features based only on ReLU activation.
@@ 4 Evaluating MSAE
In this section, we conduct extensive experiments to evaluate MSAE against ReLU and TopK SAEs. We compare the sparsity–fidelity trade-off (Section 4.2), at multiple granularity levels  (Section 4.3). We follow with evaluating the semantic quality of learned representations beyond traditional distance metrics (Section 4.4), analyzing decoder orthogonality (Section 4.5), and examining the statistical properties of SAE activation magnitudes (Section 4.6). To verify that MSAE successfully learns hierarchical features, we conduct experiments on the progressive recovery task (Section 4.7). We conclude with an ablation study comparing the influence of different training modalities in Section 4.8. In response to reviewer feedback, we’ve incorporated an analysis of BatchTopK models within Section 4.2. However, given their performance characteristics were not as competitive, we didn’t extend their evaluation to subsequent sections.
@@
Table 1: Quantitative comparison of SAE models on ImageNet-1k. We compare the following SAEs with expansion rate 8: ReLU with varying sparsity regularization ($\lambda$), TopK and BatchTopK with 64 or 256 active neurons, and Matryoshka using uniform (UW) or reverse weighting (RW) $\alpha$ coefficients. Arrows ($\uparrow$/$\downarrow$) indicate the preferred direction of metrics. NDN values in parentheses show the dead neuron count on the training set. LP (KL) values are scaled by $10^{6}$ for readability. Extended results for higher expansion rates and the text modality are reported in Appendix F.5.
@@ 4.1 Evaluation Metrics
$L_{0}$ denotes the mean proportion of zero elements in SAE activations.
Fraction of variance unexplained (FVU), also known as Normalized MSE (Gao et al., 2025), measures reconstruction fidelity by normalizing the mean squared reconstruction error $\mathcal{L}(x)$ by the mean squared value of the (mean-centered) input. Explained variance ratio (EVR) is FVU’s complement metric, defined as $1-\text{FVU}$.
Linear probing (LP) assesses how well SAE preserves semantic information in the reconstructed embeddings on the downstream task. To evaluate this, we train a linear probe model on ImageNet-1k using CLIP embeddings as a backbone, with AdamW optimizer ($\mathrm{lr}=1e{-}3$), ReduceLROnPlateau scheduler, and batch_size=256. We measure performance by comparing predictions from original versus reconstructed embeddings using two metrics: Kullback-Leibler divergence (KL) between predicted class distributions and classification accuracy (Acc), where accuracy uses $\mathrm{argmax}$ predictions from original embeddings as targets.
Centered kernel nearest neighbor alignment (CKNNA) (Huh et al., 2024) measures kernel alignment based on mutual nearest neighbors, providing a quantitative assessment of alignment between SAE activations and input embeddings. A detailed explanation is provided in Appendix E.
Decoder orthogonality (DO) calculates the mean cosine similarity of the lower triangular portion of the SAE decoder, where 0 indicates perfect orthogonality. This metric assesses how orthogonal the monosemantic feature directions are in the decoder.
Number of dead neurons (NDN) is a metric that measures how many neurons remain consistently inactive (zero in the SAE activations layer) across all inputs during training or evaluation, indicating the network’s inability to fully utilize its capacity for learning semantic features.
Figure 2: Comparison of sparsity–fidelity trade-offs across SAE architectures on ImageNet-1k. Each model presents results from all 3 expansion rates, comparing ReLU SAE ($\lambda=\{0.03,0.01,0.003\}$), TopK SAE ($k=64,128,256\}$), BatchTopK SAE ($k=64,128,256\}$) and MSAE (RW, UW). The optimal SAE would occupy the upper right corner, achieving both high sparsity and reconstruction fidelity. For extended results across both modalities, refer to Figure 11.
@@ 4.2 Sparsity–Fidelity Trade-off
We assess SAE performance using sparsity–fidelity trade-off, measuring sparsity with $L_{0}$ and reconstruction quality with EVR, following previous work. Figure 2 reveals that ReLU SAE shows difficulty balancing performance, achieving either high fidelity with low $L_{0}$ or the opposite, with expansion rate primarily improving sparsity. TopK SAE with higher $k$ values achieves better but not ReLU-level fidelity while offering improved sparsity yet consistently suffering from at least 5% of dead neurons (Table 1). While the BatchTopK variant performs similarly to TopK, it exhibits higher fidelity with lesser sparsity when trained on the same $k$. Both variants of MSAE achieve better sparsity than ReLU and better fidelity than TopK or BatchTopK, establishing a superior Pareto frontier while maintaining less than 1% of dead neurons. The RW variant further improves sparsity as expected, with only minor fidelity degradation. Notably, only Matryoshka consistently improves on both metrics with higher expansion rates, while TopK struggles with reconstruction, BatchTopK with sparsity, and ReLU shows improvements only in the highest $\lambda$ at increased expansion rates.